a long distance phone call costs $0.25 for the first minute and $0.10 for each additional minute or portion thereof. Write an inequality representing the number of minutes a person could talk without exceeding $5.

Grph the solution set of x>5.

Answer by ptaylor(2198)

Let x=number of minutes a person could talk without exceeding $5
Cost of first minute = $0.25
Cost of remaining minutes = $0.10(x-1)---must subtract out first minute
So, our inequality to solve is:
$0.25+$0.10(x-1) is less than or equal to (<=)$5.00 -----------------ans

Lets go ahead and solve the inequality:
First we'll get rid of parens
$0.25+0.10x-$0.10<=$5.00 combine like terms
$0.15+$0.10x<=$5.00 subtract $0.15 from both sides
$0.10x<=$4.85 divide both sides by 0.10
x<=48.5 minutes-------------------round down to 48 min (why? because the cost is for each additional minute or portion thereof and if we rounded up, the inequality would no longer be less than or equal to 48.5)
CK
1st minute =$0.25
47 remaining minutes=($0.10*47)=$4.70
$4.70+0.25=$4.95 note that one more minute would break the bank

Hope this helps---ptaylor